Calculate the bond equivalent yield and effective annual return on fed funds that are 10 days from maturity and have a quoted yield of 0.25 percent. (Use 365 days in a year. Do not round intermediate calculations. Round your answers to 4 decimal places. (e.g., 32.1616))
Answer #1

Bond equivalent yield = 0.25%(365/360) = 0.2535%

Effective annual return = (1 + (0.002535)/(365/10))365/10 - 1 = 0.2538%